The angular acceleration of an rotating body is define by the relation 𝔞 = - kθ. Determine a)

the value of k if angular velocity is 8 rad/sec at angular displacement of 0 radian and angular

velocity is 0 rad/sec at angular displacement of 4 radian. B) find the angular displacement at

angular displacement of 3 radian.


1
Expert's answer
2021-05-05T07:40:54-0400

"w² = w_0² +2\alpha \theta"


"a. \ 8² = w_0² + 2\alpha (0)\\
\quad64 = w_0² + 0\\
\quad w_0 = 8\ rad/s"


"0² = w_0² + 2\alpha(4)\\
0 = 8² + 8\alpha\\
\alpha = \dfrac{-64}{8} = -8\ rad/s²"



b. "w² = 8² + 2(-8)(3)\\"

"\quad w² = 64 -48\\
\quad w= \sqrt{16} = 4\ rad/s"